OpenStudy (gamereternity):

Raisins sell for $3.50 per pound, and granola sells for $5.90 per pound. Terri bought some raisins and some granola. The total weight was 2.1 pounds and cost $8.79. Let r represent the number of pounds of raisins Terri purchased. Which expression represents the number of pounds of granola she purchased? A. 2.1 – r<=my answer B. 2.1r C. r + 2.1 D. r – 2.1

OpenStudy (gamereternity):

@Mehek14

OpenStudy (gamereternity):

@Awolflover1

OpenStudy (gamereternity):

@jim_thompson5910 @kittiwitti1

jimthompson5910 (jim_thompson5910):

r = number of pounds of raisins g =number of pounds of granola `The total weight was 2.1 pounds` so r+g = 2.1 Now solve for g. To do that, subtract r from both sides r+g = 2.1 r+g-r = 2.1-r g = 2.1-r So you have the correct answer

OpenStudy (gamereternity):

ty very much

jimthompson5910 (jim_thompson5910):

no problem